# On Edge Coloring of Multigraphs

### v4

@@ 6 Conclusion
Following HochNS80 by Hochbaum et al. and StiebSTF-Book by Stiebitz et al., we may always assume that $k=O(\Delta)$, where $\Delta=\Delta(G)$ is the maximum degree of $G$. Then, each set $\overline{\varphi}(x)$ can be found in $O(\Delta)$, and therefore, one can decided in time $O(\Delta)$ whether two vertices have a common missing color. More importantly, the coloring $\varphi/P$ or $\varphi/(G-T,\alpha,\beta)$ can be constructed in time $O(|V(G)|+\Delta)$. In the proof of (A1), we used two nested “minimal” assumptions (first $p(T)$, the path number and then $|T-T_{n,q}|$). In the general proof, we also used two parameters $n,q$, where $n$ is the rung number and $q$ is the level of the hierarchy. Note that the size of an elementary set can not exceed $\Delta$ and $|V(G)|$. Thus, the four parameters $p(T)$, $|T-T_{n,q}|$, $n$ and $q$ are all bounded by $\min\{\Delta,|V(G)|\}$. Our proof is actually an algorithm to find a smaller non-elementary ETT by reducing the lexicographic order of $n\times q\times p(T)\times|T-T_{n,q}|$ through Kempe changes of types $\varphi/P$ and $\varphi/(G-T,\alpha,\beta)$. So it is not difficult to see that the complexity of our coloring algorithm is $O(|E||V|\Delta^{5})$, where $E=E(G)$ and $V=V(G)$. Moreover, if $k$ preconfigured by calculating $\max\{\Gamma(G),\Delta+1\}$, the coloring process can be done in $O(|E||V|\Delta^{4})$.
For instant future works, with this new proof of the Goldberg-Seymour conjecture, we now can start to find algorithms for all previous results based on the conjecture itself or the proof of the conjecture. For example, Cao, Chen, Ding, Zang, and the author gupta obtained some partial results towards Gupta’s co-density conjecture, which is commonly known as a dual version of the Goldberg-Seymour conjecture on edge cover packing. This new approach could be used to develop a polynomial-time algorithm for finding the corresponding edge covers. In the Ph.D Dissertation of Hao fchromatic , the authors improved the proof of the Goldberg-Seymour conjecture to confirm a generalized version of the Goldberg-Seymour conjecture on the $f$-chromatic index, see Chapter $8$ of book StiebSTF-Book for more details on the $f$-chromatic index conjecture. It might also be possible that this new approach could be revised for a polynomial-time algorithm for $f$-chromatic index.

### v5

@@ 6 Conclusion
Following [10] by Hochbaum et al. and [13] by Stiebitz et al., we may always assume that $k=O(\Delta)$, where $\Delta=\Delta(G)$ is the maximum degree of $G$. Then, each set $\overline{\varphi}(x)$ can be found in $O(\Delta)$, and therefore, one can decided in time $O(\Delta)$ whether two vertices have a common missing color. More importantly, the coloring $\varphi/P$ or $\varphi/(G-T,\alpha,\beta)$ can be constructed in time $O(|V(G)|+\Delta)$. In the proof of (A1), we used two nested “minimal” assumptions (first $p(T)$, the path number and then $|T-T_{n,q}|$). In the general proof, we also used two parameters $n,q$, where $n$ is the rung number and $q$ is the level of the hierarchy. Note that the size of an elementary set can not exceed $\Delta$ and $|V(G)|$. Thus, the four parameters $p(T)$, $|T-T_{n,q}|$, $n$ and $q$ are all bounded by $\min\{\Delta,|V(G)|\}$. Our proof is actually an algorithm to find a smaller non-elementary ETT by reducing the lexicographic order of $n\times q\times p(T)\times|T-T_{n,q}|$ through Kempe changes of types $\varphi/P$ and $\varphi/(G-T,\alpha,\beta)$. So it is not difficult to see that the complexity of our coloring algorithm is $O(|E|(|V|+\Delta)\Delta^{6})$, where $E=E(G)$ and $V=V(G)$. Moreover, if $k$ is preconfigured by calculating $\max\{\Gamma(G),\Delta+1\}$, the coloring process can be done in $O(|E|(|V|+\Delta)\Delta^{5})$.
For instant future works, with this new proof of the Goldberg-Seymour conjecture, we now can start to find algorithms for all previous results based on the conjecture itself or the proof of the conjecture. For example, Cao, Chen, Ding, Zang, and the author [3] obtained some partial results towards Gupta’s co-density conjecture, which is commonly known as a dual version of the Goldberg-Seymour conjecture on edge cover packing. This new approach could be used to develop a polynomial-time algorithm for finding the corresponding edge covers. In the Ph.D Dissertation of Hao [9], the authors improved the proof of the Goldberg-Seymour conjecture to confirm a generalized version of the Goldberg-Seymour conjecture on the $f$-chromatic index, see Chapter $8$ of book [13] for more details on the $f$-chromatic index conjecture. It might also be possible that this new approach could be revised for a polynomial-time algorithm for $f$-chromatic index.